OpenStudy (anonymous):

A rectangular swimming pool has a perimeter of 96 ft. The area of the pool is 504 ft2. Which system of equations models this situation correctly?

OpenStudy (abmon98):

P=2(l+w) A=L*W 96=2l+2W 504=LW